# CECILIA: Gas-Phase Physical Conditions and Multi-Element Chemistry at Cosmic Noon

### v1

@@ 4.2 S/O and Ar/O
These direct abundance patterns complement previous evidence for predominant enrichment from CCSNe during Cosmic Noon. For example, jointly modeling the emission line properties and shape of the UV continuum in SFGs at Cosmic Noon has led to the conclusion that these galaxies have a combination of Fe-poor stars and modest gas-phase O enrichment (C. C. Steidel et al., 2016; A. L. Strom et al., 2018; F. Cullen et al., 2019, 2021; T. M. Stanton et al., 2024). This O/Fe enhancement is another signature of significant CCSNe enrichment, which results in O/Fe in excess of the solar ratio (K. Nomoto et al., 2006). However, measuring this O/Fe enhancement in the gas-phase is exceptionally challenging owing to the faint lines required, significant dust depletion of Fe, and fluorescent effects of the optical Fe CELs (M. Rodríguez, 1999; M. Rodríguez & R. H. Rubin, 2005; J. E. Méndez-Delgado et al., 2024). The magnitude of O/Fe enhancement is often inferred through the use of photoionization modeling (e.g., A. L. Strom et al., 2018; R. L. Sanders et al., 2020) or through simultaneous rest-UV and optical observations (F. Cullen et al., 2021; T. M. Stanton et al., 2024), and these approaches indicate a O/Fe enhancement of 2.2(O/Fe)⊙ in galaxies at z2 (F. Cullen et al., 2019; A. L. Strom et al., 2022). The weighted average O/Ar observed in the combined CECILIA and literature sample at 2z3.1 (23 galaxies total) is 1.730.15(O/Ar)⊙, similar to the findings of T. M. Stanton et al. (2025).
@@ 4.3 N/O
The final abundance ratio we discuss in the CECILIA galaxies is N/O, which we plot against the total O/H abundance in Figure 10. We focus our comparison relative to the local star-forming nebulae and EELGs, and we have also included the N/O-O/H evolution observed in Galactic stars (D. C. Nicholls et al., 2017). Among the abundance ratios discussed in this work, N/O is the most sensitive to the relative flux calibration between the MOSFIRE and NIRSpec data. This is because N/O is measured entirely from [N II] and [O II] emission lines and the low-ionization zone Te; as such, any systematic uncertainty that affects Te[O II] (e.g., reddening, MOSFIRE slit-loss corrections, comparison of [O II]-to-H I intensities) will directly influence N+, O+, and the N/O relative abundance.
@@
In general, the CECILIA galaxies are consistent with the N/O-O/H trends of local SFGs, albeit with large uncertainty and significant scatter. This is similar to the N/O-O/H trends in the full KBSS sample measured via photoionization modeling (see Figure 16 in A. L. Strom et al., 2018). At z0, there are two distinct sequences in the N/O-O/H: at 12+log(O/H)8.0, the relative abundance of N/O remains constant at the ”primary plateau” of 1.450.10 dex (D. R. Garnett, 1990; L. van Zee & M. P. Haynes, 2006), which has been attributed to metallicity-insensitive N and O enrichment from massive stars. As O/H increases, metallicity-sensitive (secondary) N production from low- and intermediate-mass stars (LIMS) overtakes the N enrichment from massive stars, resulting in an increase in N/O after a sufficient time delay (R. B. C. Henry et al., 2000; F. Vincenzo et al., 2016).
@@
The large scatter in the CECILIA data may suggest that some high-z galaxies occupy two intriguing areas of the abundance ratio diagram. First, the most metal-rich galaxies have log(N/O) only slightly above the z0 primary plateau. Second, some lower-metallicity CECILIA galaxies show log(N/O) 1.1 dex, more than 0.3 dex above the primary plateau. These abundance patterns are unexpected if a galaxy has had sufficient time for significant secondary enrichment from the LIMS. However, SFE plays a significant role in the position of a galaxy in the N/O-O/H diagram. For example, galaxies with high SFE can rapidly enrich with O before the N/O ratio responds to the N enrichment from LIMS. In this way, log(N/O) ratios near the primary plateau can be achieved at O/H in excess of the solar abundance ratio (R. B. C. Henry et al., 2000; F. Vincenzo et al., 2016). Similarly, galaxies with low SFE maintain low O/H while the N abundance gradually increases through LIMS enrichment, producing high N/O at low O/H. Other mechanisms can also produce elevated N/O at low metallicity, such as local pollution from Wolf-Rayet stars (e.g., B. L. James et al., 2009) or significant pristine gas infall into a moderately N-enriched ISM.
@@
Recent JWST observations suggest that elevated N/O at low O/H may be common in high-z SFGs. Observations of galaxies at z6 have revealed the surprising detection of UV N III and N IV emission lines, which are rarely observed in local nebulae (e.g., M. Mingozzi et al., 2022). The N/O abundances implied from the UV emission lines in these high-z EELGs are in excess of 0.75 dex at 12+log(O/H)8.0 dex (P. Senchyna et al., 2024; R. Marques-Chaves et al., 2024; M. W. Topping et al., 2024, 2025b; M. Curti et al., 2025). While UV N/O may be affected by systematic uncertainties, such as the ICF and density effects, the N/O ratios measured from the optical [N II] and [O II] emission lines in other high-z galaxies also indicate an enhancement in N with respect to O at moderate metallicities (L. Clarke et al., 2023; R. L. Sanders et al., 2023a; B. Welch et al., 2024, 2025; T. Morishita et al., 2025). Recently, K. Z. Arellano-Córdova et al. (2025) measured the N/O-O/H abundance patterns in two SFGs in the EXCELS survey and compiled all N/O abundances in galaxies at z1.8, including those measured from rest-optical and UV emission lines. All galaxies were found to have 12+log(O/H)8.2 dex and log(N/O)1.10 dex, except D40 which has a N/O-O/H abundance pattern consistent with the primary plateau.
@@
The diversity of N/O-O/H abundances measured in the CECILIA galaxies may provide additional evidence for variations in this abundance ratio beyond the expectations of local SFGs. If the large scatter in N/O-O/H diagram is observed in other high-z galaxies, then the differences in the chemical abundance ratios from the local nebulae have significant implications for photoionization modeling: many model grids assume a range of N/O that mirrors the N/O-O/H relation derived from local nebulae. Given the sensitivity of N/O on the SFH and SFEs, it may be necessary to fully decouple N/O and O/H when modeling the spectra of high-z galaxies (e.g., N. Vale Asari et al., 2016; A. L. Strom et al., 2018). Chemical evolution models have been able to match the elevated N/O observed in some high-z EELGs by invoking bursty SFHs with varying SFE (e.g., C. Kobayashi & A. Ferrara, 2024). Indeed, differences in SFH and SFE can manifest as different locations of the N/O plateau, a trend that has been measured from resolved H II regions within local galaxies (see F. Belfiore et al., 2017; D. A. Berg et al., 2020). Further constraints on the gas-phase N/O during Cosmic Noon will reveal the prevalence of the N-enhanced ISM and whether these trends require bursty SFHs. Additionally, comparisons between N/O abundances simultaneously measured from joint UV and optical spectroscopy (in high-z and local SFGs) are required to understand the mechanisms of UV N emission and whether these independent measurements are representative of the average N/O abundance in the ISM of high-z SFGs.
@@ 5 Conclusions
In this paper, we have presented the gas-phase physical conditions and chemical abundances in CECILIA, a Cycle 1 JWST program (PID2593) targeting SFGs at Cosmic Noon to unveil the buildup of heavy elements during the peak of cosmic star formation. CECILIA combines ultra-deep JWST/NIRSpec observations (29.5hr in G235M, 1.1hr in G395M) with photometric and spectroscopic data of each target galaxy acquired from KBSS, enabling a detailed examination of the conditions in the multi-phase ISM. Specifically, the faint [S III] and [O II] auroral lines are detected, allowing for the direct determination of Te and the chemical abundance patterns from the strong rest-optical and NIR metal CELs in 20 SFGs at Cosmic Noon. The combination of these new NIRSpec data with the archival MOSFIRE observations results in one of the largest samples of direct-method abundances at z2 to date. The main aspects of our investigation are:
@@
We measure the electron density in the ISM via the [S II]6718,33 doublet in the NIRSpec data and the [O II]3727,29 lines in the higher-resolution MOSFIRE spectra (§3.1). A simultaneous measurement of ne[S II] and ne[O II] is possible in 15 CECILIA galaxies (Figure 4); the average densities measured in these 15 galaxies are ne[S II]26744 and ne[O II]39573 cm-3, approximately a factor of two larger than the typical ne measured in local SFGs. While the average ne[O II] is slightly larger than ne[S II], the similar ne in the S+ and O+ emitting gas may be consistent with minimal DIG contamination to the [S II] line ratios or higher ne in the DIG of galaxies at Cosmic Noon. There are galaxies in the CECILIA sample with ne[S II]100 cm-3 (top panel of Figure 4), where assuming a characteristic ne 250 cm-3 is not appropriate for ionic abundance calculations.
@@
Using the ne and Te measured in the CECILIA galaxies, we calculate the N, O, S, and Ar abundances in the ISM. The O/H abundances measured in the CECILIA galaxies range from 12+log(O/H) 7.76 to 8.78 dex (left panel of Figure 6). The CECILIA sample contains the highest direct-method O/H abundances measured with JWST to date, indicating that galaxies can reach solar O/H by Cosmic Noon. The O/H abundances are in qualitative agreement with the strong-line trends observed in the same galaxies (right panel of Figure 6). We discuss potential sources of systematic uncertainties, noting that the Te measured in the CECILIA galaxies are more akin to metal-rich H II regions and that the same abundance results are recovered if Te[S III] (which is measured entirely from NIRSpec) is used to measure O/H.
@@
The S/O and Ar/O abundances measured in the CECILIA galaxies show no evolution with O/H, but they are significantly sub-solar (Figure 8). This trend is inconsistent with the chemical abundance patterns of many local SFGs, which mostly exhibit solar S/O and Ar/O. However, both S and Ar have a non-negligible Type Ia enrichment component (C. Kobayashi et al., 2020) such that S/O and Ar/O are sensitive to the delay between CCSNe and Type Ia enrichment. The low S/O and Ar/O, therefore, suggests that the typical galaxy at Cosmic Noon has been predominantly enriched with the nucleosynthetic products of CCSNe. This trend is consistent with prior inferences of the Fe/O abundance in similar galaxy populations. We examine the redshift evolution of the Ar/O ratio using CECILIA and a compilation of direct Ar/O measurements from the literature, finding that the average Ar/O at z2-2.5 is in good agreement with IMF-averaged CCSNe yield predictions for metal-poor stars (Figure 9).
@@ Appendix A Global Background Subtraction
It is reassuring that both methods produce relatively consistent strong line fluxes; indeed, other studies have found consistent line fluxes when using msaexp’s recently-implemented global background strategy and the default nod background approach (e.g., A. de Graaff et al., 2024b, 2025). This trend is related to the optimal extraction profiles, which provide the most weight to the central 1-2 pixels in the spatial direction when producing the 1D spectra. We do note that there is relatively large scatter in the observed line flux ratios, where the global background approach produces generally stronger [S II]6718. The change in emission line flux is related to the chosen background approach and the resulting optimal 1D extraction, which can weight different portions of the 2D spectra based on the assumed profile. In summary, the global background approach preserves the extended strong line emission while producing line fluxes similar to the nod background method when using an optimal extraction algorithm. For future spatial analyses of the CECILIA galaxies and to capture the total emission from lines such as [S II] and H, we adopt the global background subtraction method for the determination of Te and chemical abundances.
@@ Appendix B Reddening Corrections and Cross Grating Uncertainties
Figure 13 plots the trends measured from the Paschen-to-H line ratios in the CECILIA sample. The left panel focuses on the inferred from the P8/H ratio. This Paschen line is most frequently detected in the overlap portion of the G235M and G395M spectra, permitting a comparison between the predicted from the different detectors. While the uncertainties on the G395M-derived are large on account of the faint P8 line and noisy continuuum, four of the seven galaxies show agreement between the independent measurements in G235M and G395M. Observational challenges associated with robustly measuring P8 may lead to discrepant . For example, the P8 line is in close proximity to [S III]9533: while both lines are resolved in the average G235M spectrum, the two lines are blended in most G395M data.
@@
Differences to the line fits resulting from the coarser wavelength sampling and noisier spectra can contribute to the resulting scatter about the one-to-one line in the left panel, so in the right panel of Figure 13 we plot measured from G235M and G395M using the average from multiple Pachen-to-H line ratios. The from G235M and G395M scatter around the one-to-one line, although measured in individual galaxies can differ depending on the available Paschen lines. This comparison assumes the shape of the N. A. Reddy et al. (2020) is valid at wavelengths longer than 8000 Å, an area where it is not well calibrated. Differences in from the different Paschen lines, therefore, could be the result of a different attenuation law shape at NIR wavelengths. Indeed, early evidence from JWST observations of high-z galaxies indicate that the attenuation curve may deviate from both the J. A. Cardelli et al. (1989) and N. A. Reddy et al. (2020) calibrations. For example, R. L. Sanders et al. (2024b) measure the attenuation curve from eleven Balmer and Paschen lines observed in a galaxy at z4.41, finding a steeper attenuation curve in the NIR relative to the N. A. Reddy et al. (2020) parameterization (see other works on the dust attenuation in high-z galaxies by N. A. Reddy et al., 2023a; V. Markov et al., 2025).
@@
Another potential systematic is a relative flux calibration error between G235M and G395M. This is most evident when comparing the line fluxes measured in the overlap region between the G235M and G395M gratings. Depending on the galaxy’s redshift, this can include the [S III] nebular lines, P8, and P9. In Figure 14, we provide a histogram distribution of the flux ratios for these emission lines. Each ratio represents the flux measured for an emission line independently in G235M and G395M after global background subtraction, optimal extraction from the 2D spectra, and Gaussian emission line fitting. We note that [S III]9533 and P89549 are often blended in the G395M data, so we compare the G235M and G395M ratio of the total [S III]P8 flux. The Fλ(G235M)/Fλ(G395M) weighted average is 0.95 with standard deviation 0.09, indicating that the emission line fluxes measured from G235M and G395M generally agree but with non-negligible scatter.
@@
These line flux variations can introduce systematic differences into the derived from the G395M Paschen lines relative to the G235M H flux. It is for this reason that we adopt the neighboring H I line approach for a pseudo-reddening correction (see Section 3.2.1), as this bypasses the uncertainties in the attenuation curve and relative flux calibration. Additionally, when reporting we prioritize measurements from the Balmer decrement in NIRSpec and MOSFIRE since the N. A. Reddy et al. (2020) attenuation curve is calibrated at these wavelengths. In instances when H and H are not significantly detected, we use constrained from the Paschen lines and H. Finally, we adopt a 9% uncertainty on all flux ratios involving lines in G235M and G395M (e.g., Paschen-to-H ratios) to account for the observed scatter in the individual line flux measurements. In the future, it may be possible to constrain the average shape of the NIR attenuation curve for galaxies at 2z3 from the CECILIA spectra, but we do not attempt this calibration in the present work owing to the persistent fluxing issues.

### v2

@@ 4.2 S/O and Ar/O
These direct abundance patterns complement previous evidence for predominant enrichment from CCSNe during Cosmic Noon. For example, jointly modeling the emission line properties and shape of the UV continuum in SFGs at Cosmic Noon has led to the conclusion that these galaxies have a combination of Fe-poor stars and modest gas-phase O enrichment (steidel2016; strom2018; cullen2019; cullen2021; stanton2024). This O/Fe enhancement is another signature of significant CCSNe enrichment and a dearth of Type Ia SNe products, which results in O/Fe in excess of the solar ratio (nomoto2006). However, measuring this O/Fe enhancement in the gas-phase is exceptionally challenging owing to the faint lines required, significant dust depletion of Fe, and fluorescent effects of the optical Fe CELs (rodriguez1999; rodriguez2005; mendez-delgado2024). The magnitude of O/Fe enhancement is often inferred through the use of photoionization modeling (e.g., strom2018; sanders2020), simultaneous rest-UV and optical observations (cullen2021; stanton2024), or direct measurement from optical CELs (curti2025b), and these approaches indicate a O/Fe enhancement of 2.2(O/Fe)⊙ in galaxies at z2 (cullen2019; strom2022). The inverse variance-weighted average O/Ar observed in the combined CECILIA and literature sample at 2z3.1 (23 galaxies total) is 1.780.16(O/Ar)⊙, similar to the findings of stanton2025.
@@ 4.3 N/O
The final abundance ratio we discuss in the CECILIA galaxies is N/O, which we plot against the total O/H abundance in Figure 10. We focus our comparison relative to the local star-forming nebulae and EELGs, and we have also included the N/O-O/H evolution observed in Galactic stars (nicholls2017). Among the abundance ratios discussed in this work, N/O is the most sensitive to the relative flux calibration between the MOSFIRE and NIRSpec data. This is because N/O is measured using only the low-ionization zone Te; as such, any systematic uncertainty that affects Te[O II] (e.g., reddening, MOSFIRE slit-loss corrections, comparison of [O II]-to-H I intensities) will directly influence N+, O+, and the N/O relative abundance.
@@
In general, the CECILIA galaxies are consistent with the N/O-O/H trends of local SFGs, albeit with large uncertainty and significant scatter. This is similar to the N/O-O/H trends in the full KBSS sample measured via photoionization modeling (see Figure 16 in strom2018). At z0, there are two distinct sequences in the N/O-O/H abundances: at 12+log(O/H)8.0, the relative abundance of N/O remains constant at the ”primary plateau” of 1.450.10 dex (garnett1990; vanZee2006), which has been attributed to metallicity-insensitive N and O enrichment from massive stars. As O/H increases, metallicity-sensitive (secondary) N production from the asymptotic giant branch phase in low- and intermediate-mass stars (LIMS) overtakes the N enrichment from massive stars, resulting in an increase in N/O after a sufficient time delay (henry2000; vincenzo2016).
@@
The large scatter in the CECILIA data may suggest that some high-z galaxies occupy two intriguing areas of the abundance ratio diagram. First, the most metal-rich galaxies have log(N/O) only slightly above the z0 primary plateau. Second, some lower-metallicity CECILIA galaxies show log(N/O) 1.1 dex, more than 0.3 dex above the primary plateau. These abundance patterns are unexpected if a galaxy has had sufficient time for significant secondary enrichment from the LIMS. However, SFE plays a significant role in the position of a galaxy in the N/O-O/H diagram. For example, galaxies with high SFE can rapidly enrich with O before the N/O ratio responds to the N enrichment from LIMS. In this way, log(N/O) ratios near the primary plateau can be achieved at O/H in excess of the solar abundance ratio (henry2000; vincenzo2016). Similarly, galaxies with low SFE maintain low O/H while the N abundance gradually increases through LIMS enrichment, producing high N/O at low O/H. Other mechanisms can also produce elevated N/O at low metallicity, such as local pollution from Wolf-Rayet stars (e.g., james2009), significant pristine gas infall into a moderately N-enriched ISM, or from bursty star formation and strong stellar feedback (e.g., mcclymont2025; berg2025).
@@
Recent JWST observations suggest that elevated N/O at low O/H may be common in high-z SFGs. Observations of galaxies at z6 have revealed the surprising detection of UV N III and N IV emission lines, which are rarely observed in local nebulae (e.g., mingozzi2022; martinez2025). The N/O abundances implied from the UV emission lines in these high-z EELGs are in excess of 0.75 dex at 12+log(O/H)8.0 dex (senchyna2024; marques-chaves2024; topping2024; topping2025no; curti2025). While UV N/O may be affected by systematic uncertainties, such as the ICF and density effects, the N/O ratios measured from the optical [N II] and [O II] emission lines in other high-z galaxies also indicate an enhancement in N with respect to O at moderate metallicities (clarke2023; sanders2023; welch2024; welch2025; morishita2025). Particularly, arellano-cordova2025 measured the N/O-O/H abundance patterns in two SFGs in the EXCELS survey and compiled all N/O abundances in galaxies at z1.8, including those measured from rest-optical and UV emission lines. All galaxies were found to have 12+log(O/H)8.2 dex and log(N/O)1.10 dex, except D40 which has a N/O-O/H abundance pattern consistent with the primary plateau.
@@
The diversity of N/O-O/H abundances measured in the CECILIA galaxies may provide additional evidence for variations in this abundance ratio beyond the expectations of local SFGs. If the large scatter in N/O-O/H diagram is observed in other high-z galaxies, then the differences in the chemical abundance ratios from the local nebulae have significant implications for photoionization modeling: many model grids assume a range of N/O that mirrors the N/O-O/H relation derived from local nebulae. Given the sensitivity of N/O on the SFH and SFEs, it may be necessary to fully decouple N/O and O/H when modeling the spectra of high-z galaxies (e.g., valeasari2016; strom2018). Chemical evolution models have been able to match the elevated N/O observed in some high-z EELGs by invoking bursty SFHs with varying SFE (e.g., kobayashi2024; mcclymont2025). Indeed, differences in SFH and SFE can manifest as different locations of the N/O plateau, a trend that has been measured from resolved H II regions within local galaxies (see belfiore2017; berg2020). Further constraints on the gas-phase N/O during Cosmic Noon will reveal the prevalence of the N-enhanced ISM and whether these trends require bursty SFHs. Additionally, comparisons between N/O abundances simultaneously measured from joint UV and optical spectroscopy (in high-z and local SFGs) are required to understand the mechanisms of UV N emission and whether these independent measurements are representative of the average N/O abundance in the ISM of high-z SFGs.
@@ 5 Conclusions
In this paper, we have presented the gas-phase physical conditions and chemical abundances in CECILIA, a Cycle 1 JWST program targeting SFGs at Cosmic Noon to unveil the buildup of heavy elements during the peak of cosmic star formation. CECILIA combines ultra-deep JWST/NIRSpec observations (29.5hr in G235M, 1.1hr in G395M) with photometric and spectroscopic data of each target galaxy acquired from KBSS, enabling a detailed examination of the conditions in the multi-phase ISM. Specifically, the faint [S III] and [O II] auroral lines are detected, allowing for the direct determination of Te and the chemical abundance patterns from the strong rest-optical and NIR metal CELs in 20 SFGs at Cosmic Noon. The combination of these new NIRSpec data with the archival MOSFIRE observations results in one of the largest samples of direct-method abundances at z2 to date. The main aspects of our investigation are:
@@
We measure the electron density in the ISM via the [S II]6718,33 doublet in the NIRSpec data and the [O II]3727,29 lines in the higher-resolution MOSFIRE spectra (§3.1). A simultaneous measurement of ne[S II] and ne[O II] is possible in 15 CECILIA galaxies (Figure 4); the average densities measured in these 15 galaxies are ne[S II]26744 and ne[O II]38670 cm-3, approximately a factor of two larger than the typical ne measured in local SFGs. While the average ne[O II] is slightly larger than ne[S II], the similar ne in the S+ and O+ emitting gas may be consistent with minimal DIG contamination to the [S II] line ratios or higher ne in the DIG of galaxies at Cosmic Noon. There are galaxies in the CECILIA sample with ne[S II]100 cm-3 (top panel of Figure 4), where assuming a characteristic ne 250 cm-3 is not appropriate for ionic abundance calculations.
@@
Using the ne and Te measured in the CECILIA galaxies, we calculate the N, O, S, and Ar abundances in the ISM. The O/H abundances measured in the CECILIA galaxies range from 12+log(O/H) 7.76 to 8.81 dex (left panel of Figure 6). The CECILIA sample contains the highest direct-method O/H abundances measured with JWST to date, indicating that galaxies can reach solar O/H by Cosmic Noon. The O/H abundances are in qualitative agreement with the strong-line trends observed in the same galaxies (right panel of Figure 6). We discuss potential sources of systematic uncertainties, noting that the Te measured in the CECILIA galaxies are more akin to metal-rich H II regions and that the same abundance results are recovered if Te[S III] (which is measured entirely from NIRSpec) is used to measure O/H.
@@
The S/O and Ar/O abundances measured in the CECILIA galaxies show no evolution with O/H, but they are significantly sub-solar (Figure 8). This trend is inconsistent with the chemical abundance patterns of many local SFGs, which mostly exhibit solar S/O and Ar/O. However, both S and Ar have a non-negligible Type Ia enrichment component (kobayashi2020SNIa) such that S/O and Ar/O are sensitive to the delay between CCSNe and Type Ia enrichment. The low S/O and Ar/O, therefore, suggests that the typical galaxy at Cosmic Noon has been predominantly enriched with the nucleosynthetic products of CCSNe. This trend is consistent with prior inferences of the Fe/O abundance in similar galaxy populations. We examine the redshift evolution of the Ar/O ratio using CECILIA and a compilation of direct Ar/O measurements from the literature, finding that the average Ar/O at z2-2.5 is in good agreement with IMF-averaged CCSNe yield predictions for metal-poor stars (Figure 9).
@@ Appendix A Global Background Subtraction
It is reassuring that both methods produce relatively consistent strong line fluxes; indeed, other studies have found consistent line fluxes when using msaexp’s recently-implemented global background strategy and the default nod background approach (e.g., degraaf2024; degraaf2025). This trend is related to the optimal extraction profiles, which provide the most weight to the central 1-2 pixels in the spatial direction when producing the 1D spectra. We do note that there is relatively large scatter in the observed line flux ratios, where the global background approach produces generally stronger [S II]6718. The change in emission line flux is related to the chosen background approach and the resulting optimal 1D extraction, which can weight different portions of the 2D spectra based on the assumed profile. In summary, the global background approach preserves the extended strong line emission while producing line fluxes similar to the nod background method when using an optimal extraction algorithm. For future spatial analyses of the CECILIA galaxies and to capture the total emission from lines such as [S II] and H, we adopt the global background subtraction method for the determination of Te and chemical abundances.
@@ Appendix B Reddening Corrections and Cross Grating Uncertainties
Figure 13 plots the trends measured from the Paschen-to-H line ratios in the CECILIA sample. The left panel focuses on the inferred from the P8/H ratio. This Paschen line is most frequently detected in the overlap portion of the G235M and G395M spectra, permitting a comparison between the predicted from the different detectors. While the uncertainties on the G395M-derived are large on account of the faint P8 line and noisy continuum, four of the seven galaxies show agreement between the independent measurements in G235M and G395M. Observational challenges associated with robustly measuring P8 may lead to discrepant . For example, the P8 line is in close proximity to [S III]9533: while both lines are resolved in the average G235M spectrum, the two lines are blended in most G395M data.
@@
Differences to the line fits resulting from the coarser wavelength sampling and noisier spectra can contribute to the resulting scatter about the one-to-one line in the left panel, so in the right panel of Figure 13 we plot measured from G235M and G395M using the average from multiple Pachen-to-H line ratios. The from G235M and G395M scatter around the one-to-one line, although measured in individual galaxies can differ depending on the available Paschen lines. This comparison assumes the shape of the reddy2020 is valid at wavelengths longer than 8000 Å, an area where it is not well calibrated. Differences in from the different Paschen lines, therefore, could be the result of a different attenuation law shape at NIR wavelengths. Indeed, early evidence from JWST observations of high-z galaxies indicate that the attenuation curve may deviate from both the cardelli1989 and reddy2020 calibrations. For example, sanders2024 measure the attenuation curve from eleven Balmer and Paschen lines observed in a galaxy at z4.41, finding a steeper attenuation curve in the NIR relative to the reddy2020 parameterization (see other works on the dust attenuation in high-z galaxies by reddy2023; markov2025).
@@
Another potential systematic is a relative flux calibration error between G235M and G395M. This is most evident when comparing the line fluxes measured in the overlap region between the G235M and G395M gratings. Depending on the galaxy’s redshift, this can include the [S III] nebular lines, P8, and P9. In Figure 14, we provide a histogram distribution of the flux ratios for these emission lines. Each ratio represents the flux measured for an emission line independently in G235M and G395M after global background subtraction, optimal extraction from the 2D spectra, and Gaussian emission line fitting. We note that [S III]9533 and P89549 are often blended in the G395M data, so we compare the G235M and G395M ratio of the total [S III]P8 flux. The Fλ(G235M)/Fλ(G395M) inverse variance weighted average is 0.95 with uncertainty of 0.09, indicating that the emission line fluxes measured from G235M and G395M generally agree but with non-negligible scatter.
@@
These line flux variations can introduce systematic differences into the derived from the G395M Paschen lines relative to the G235M H flux. It is for this reason that we adopt the neighboring H I line approach for a pseudo-reddening correction (see Section 3.2.1), as this bypasses the uncertainties in the attenuation curve and relative flux calibration. Additionally, when reporting we prioritize measurements from the Balmer decrement in NIRSpec and MOSFIRE since the reddy2020 attenuation curve is calibrated at these wavelengths. In instances when H and H are not significantly detected, we use constrained from the Paschen lines and H. Finally, we adopt a 9% uncertainty on all flux ratios involving lines in G235M and G395M (e.g., Paschen-to-H ratios) to account for the observed scatter in the individual line flux measurements. In the future, it may be possible to constrain the average shape of the NIR attenuation curve for galaxies at 2z3 from the CECILIA spectra, but we do not attempt this calibration in the present work owing to the persistent fluxing issues.